OpenStudy (anonymous):

Solve y over negative 6 + 5 = 9

OpenStudy (calculusxy):

Do you mean: \[\frac{ y }{ -(6+5) } = 11\]

OpenStudy (calculusxy):

It's unclear as to what equation you need help solving for.

OpenStudy (rhr12):

Y is -11 as i think

OpenStudy (anonymous):

\[\frac{ y }{ -6 }+5=9\]

OpenStudy (rhr12):

Ow. That's different. Hold on.

OpenStudy (anonymous):

lol k

OpenStudy (rhr12):

-24 is the value of y

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

can u help me with another problem please

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }( 2x -14)=4\]

OpenStudy (rhr12):

I think its 36.

OpenStudy (rhr12):

Check it with the calculator

OpenStudy (anonymous):

no i got it its 15

OpenStudy (anonymous):

but thanks